# Tell Me What You Don’t Know: Enhancing Refusal Capabilities of Role-Playing Agents via Representation Space Analysis and Editing

### v1

@@ Abstract
Role-Playing Agents (RPAs) have shown remarkable performance in various applications, yet they often struggle to recognize and appropriately respond to hard queries that conflict with their role-play knowledge. To investigate RPAs’ performance when faced with different types of conflicting requests, we develop an evaluation benchmark that includes contextual knowledge conflicting requests, parametric knowledge conflicting requests, and non-conflicting requests to assess RPAs’ ability to identify conflicts and refuse to answer appropriately without over-refusing. Through extensive evaluation, we find that most RPAs behave significant performance gaps toward different conflict requests. To elucidate the reasons, we conduct an in-depth representation-level analysis of RPAs under various conflict scenarios. Our findings reveal the existence of rejection regions and direct response regions within the model’s forwarding representation, and thus influence the RPA’s final response behavior. Therefore, we introduce a lightweight representation editing approach that conveniently shifts conflicting requests to the rejection region, thereby enhancing the model’s refusal accuracy. The experimental results validate the effectiveness of our editing method, improving RPAs’ refusal ability of conflicting requests while maintaining their general role-playing capabilities.
@@ 1 Introduction
Role-Playing Agents(RPAs), ranging from non-player characters in video games(Wang et al., 2023a) to virtual assistants(Tseng et al., 2024) and interactive educational tools(Wei et al., 2024), are revolutionizing human-computer interaction(Chen et al., 2024b). The growing importance of RPAs in AI applications underscores the need to improve their performance. Previous work in the field of role-playing has primarily focused on enhancing the performance of RPAs through techniques such as prompt-based methods and fine-tuning(Wang et al., 2023c; Zhou et al., 2023; Tu et al., 2023; Li et al., 2023; Chen et al., 2024b; Xu et al., 2024c). To assess these improvements, researchers have introduced several fine-grained evaluation dimensions(Wang et al., 2023b; Chen et al., 2024b; d; Tu et al., 2024; Yuan et al., 2024; Tang et al., 2024; Sadeq et al., 2024), such as assess personality(Wang et al., 2023b) or hallucination(Ahn et al., 2024) of RPAs.
Although these efforts have effectively enhanced the performance of RPAs in terms of role consistency and dialogue capabilities(Wang et al., 2023c; Chen et al., 2023), RPAs often struggle when faced with queries that conflict with their role knowledge or capabilities. As a result, they tend to respond directly to queries instead of refusing to answer when faced with such conflicts(Ahn et al., 2024; Sadeq et al., 2024; Tang et al., 2024). For instance, when interacting with an RPA playing the role of Gandalf, if a user queries, “Who murdered Harry Potter’s parents?”, an ideal response would be, “I don’t know what you’re talking about. The story of Harry Potter is not part of my world or knowledge.” Instead, the RPA might incorrectly reply, “Harry Potter’s parents, James and Lily Potter, were murdered by…” Enhancing the refusal capability of RPAs is crucial for building reliable AI systems. Although some studies have begun to address this issue(Ahn et al., 2024; Sadeq et al., 2024), their scope remains limited, often focusing on specific scenarios such as temporal inconsistencies. There is a lack of systematic research on diverse conflicting scenarios and little exploration of the reasons for RPAs’ performance gap across different types of conflicting queries.
@@
Figure 1: Design of refusal scenarios. Since the knowledge basis for RPAs’ responses typically originates from contextual knowledge and parametric knowledge, we have subdivided the knowledge conflict scenarios into four categories. Among these, the role setting conflict query and role profile conflict query involve conflicts with contextual knowledge, while the factual knowledge conflict query and absent knowledge conflict query involve conflicts with the model’s parametric knowledge. Non-conflict query is used to assess the RPAs’ general role-playing ability. By analyzing the representation of these queries within the model, we find that there are rejection regions and direct response regions in the representation space. The proximity of a query to these regions largely determines the RPA’s response (rejection or direct answer).
@@
To answer RQ1 and lay the groundwork for RQ2 and RQ3, we first categorized refusal scenarios into two main categories: (1) conflicts with role contextual knowledge, and (2) conflicts with role parametric knowledge(Xu et al., 2024b). These categories were further subdivided into four specific scenarios, as illustrated in Figure 1. The expected responses from RPAs in these scenarios can range from direct refusal to acknowledging their inability to answer or providing disclaimers about potential errors. To evaluate RPAs’ refusal capabilities, we constructed an evaluation benchmark with queries designed to test various conflict scenarios. We also included non-conflicting queries to assess whether RPAs would excessively refuse to answer. Our evaluation of state-of-the-art models, including GPT-4 and Llama-3, revealed significant differences in their abilities to identify conflicts and refuse to answer across different scenarios. Notably, even advanced models showed unsatisfactory performance when dealing with queries conflicting with role parametric knowledge.
@@
To understand these performance gap, we analyzed model representations under different conflict scenarios(Zou et al., 2023; Liu et al., 2023; Li et al., 2024; Wu et al., 2024). This analysis revealed the existence of rejection regions and direct response regions within the model’s representation space. Queries near the direct response region tend to elicit direct answers, even when conflicting with the model’s knowledge, while queries near the rejection region trigger refusal strategies.
@@
Based on these findings, we developed a representation editing method to shift conflicting queries from the direct response region toward the rejection region. This approach effectively enhanced the model’s rejection capability while maintaining its general role-playing abilities. We compared our method with prompt-based and fine-tuning approaches(Wang et al., 2023c; Zhou et al., 2023; Chen et al., 2023; Li et al., 2023), demonstrating its effectiveness in rejecting conflicting queries without compromising overall performance.
2 Related Work
2.1 Role-Playing Agents
Enhancing Role-Playing Performance. Methods to improve RPAs are broadly categorized into prompt-based and fine-tuning-based approaches. Prompt-based methods provide models with detailed character descriptions, outlining attributes such as age, personality, and abilities, to facilitate accurate role-playing (Wang et al., 2023c; Zhou et al., 2023). Fine-tuning-based methods involve training models on role-specific behaviors, often using data sourced from manual annotations (Zhou et al., 2023; Chen et al., 2023; Zhang et al., 2024b), online resources(Zheng et al., 2019; Qian et al., 2021; Song et al., 2020; Shao et al., 2023; Tu et al., 2024), or generated by LLMs (Wang et al., 2023c; Li et al., 2023; Zhao et al., 2023a; Ahn et al., 2024; Lu et al., 2024). These methods aim to instill role-consistent behaviors and dialogue patterns in the models.
2.2 Knowledge Boundaries and Refusal Strategies
Parameteric Knowledge. Yang et al. (2023) and Cheng et al. (2024) explore teaching models to express uncertainty using prompt-based, fine-tuning, and preference-aware optimization methods. Xu et al. (2024a) propose a reinforcement learning method based on knowledge feedback to dynamically determine the model’s knowledge boundaries. Similarly, Zhang et al. (2024a) identifies knowledge gaps between pre-trained parameters and instruction-tuning data, constructing refusal-aware data by appending uncertainty expressions and improving the model’s ability to answer known questions while refusing unknown ones. Chen et al. (2024c) detect the knowledge boundaries of LLMs through internal confidence and teach LLMs to recognize and express these boundaries. Zhao et al. (2023b) propose a self-detection scheme to identify unknown knowledge by examining behavioral differences under varying formulations and the atypicality of input expressions. To address factual errors and outdated knowledge in parameterized knowledge, mainstream methods convert parameterized knowledge into contextual knowledge.
Contextual Knowledge. Cao (2023) use an independent structured knowledge base to represent the knowledge scope of LLMs, making LLMs process input-output data without relying on internal knowledge, thereby avoiding misinformation. Prompting LLMs to refuse to answer difficult questions improves system reliability. Deng et al. (2024) generate extensive unknown question-response data through class-aware self-augmentation and select qualified data via differential-driven self-curation, fine-tuning LLMs to improve their response capabilities to various unknown questions, enabling the model to refuse and explain why it cannot answer. Brahman et al. (2024) categorize scenarios requiring refusal to answer, and explore different training strategies to teach models to say “no.” Zhao et al. (2024) investigate decision boundaries in in-context learning by analyzing decision boundaries in binary classification tasks.
@@ 3 RoleRef: A Benchmark for Evaluating RPA’s Refusal Ability
We first introduce the scenarios where RPAs should refuse to answer. Then, based on the scenarios requiring refusal, we construct our dataset RoleRef (Role-playing agents Refuse to answer). Finally, we propose an evaluation framework to comprehensively measure the role-playing capabilities of RPAs, with a particular emphasis on how they refuse inappropriate or irrelevant questions.
@@ 3.1 Scenario Design
  • Role Profile Conflict: The user’s query is in accordance with the role profile, however, it violates specific content within the role profile. For instance, when interacting with an RPA whose role profile states “While Gandalf is powerful, he is not omnipotent.” the user asks: “Gandalf, how can I become as omnipotent as you?”
@@
  • Role’s Absent Knowledge Conflict: The character was not present when a specific event occurred. For example, when interacting with an RPA playing the role of Gandalf, the user asks: “Were you there at the moment when Goldberry bid farewell to Frodo and his friends, blessing their journey as they departed on their ponies?”.
@@ 3.2 Data Construction
For generating non-conflict queries, we directly prompt the model to generate queries and responses based on the summary content. Additionally, for each query, we require the model to provide the corresponding reference information. The prompts we used are shown in Appendix B.
  Non-conflict Role Setting Role Profile Factual Knowledge Absent Knowledge
TimeChara $6028$ $818$ $2056$
RoleRef $11838$ $16455$ $2177$ $12189$ $2104$
@@ 4.1 Models and Metrics
We evaluated these models using the RoleRef dataset. Performance was assessed across 9 dimensions (detailed in Appendix A), with GPT-4o serving as the scoring model. Each dimension was scored on a scale of 0 to 2, with the average score reported unless otherwise specified.
@@ 5.1 Analysis via Linear Probes
Previous work has shown that the internal states of LLMs can reveal the model’s knowledge about query truthfulness (Azaria & Mitchell, 2023; Ji et al., 2024). Building on this, we used linear probes to investigate whether models can distinguish between queries that should be refused and those that should be answered. The detailed procedure of probe training is provided in Appendix C.2. The results, shown in Figure 2, reveal following insight:
@@ 5.2 Analysis via t-SNE
Overlap in parametric knowledge conflicts - Direct response region. Representations of most parametric knowledge conflict queries significantly overlap with non-conflict queries. This overlap suggests that these queries within the representation space are positioned in a direct response region, where the model tends to answer directly without recognizing the conflict. For example, when presented with the query “Gandalf, was it you who recommended The Prancing Pony as a safe place to stay for Frodo and his friends?”. The representation of this query likely falls within the direct response region, leading to an inappropriate answer. Conversely, for queries whose representations fall further from the non-conflict cluster, the model correctly identifies the false and refuses to answer.
Figure 4: Methods to improve the model’s ability to refuse to answer.
6.1 Representation Editing Method
Step 1: Collecting activation. For each role, we first construct a batch of conflicting queries and non-conflicting queries. For conflicting queries, we use queries with conflicting roles. When the model responds to such conflicting queries, it can accurately identify the conflict and make corresponding rejections. When the model responds to non-conflicting queries, the rejection mechanism is often not triggered. Therefore, in both cases, we collect the internal representation of the last token position for each query, capturing both the model’s rejection and non-rejection states.
Step 2: Identifying the rejection direction. Using the collected representations, we compute the difference between conflicting and non-conflicting query representations to isolate the rejection-related features. We then calculate the cluster center of these difference vectors, termed as “rejection direction.” To refine this direction, we compute the variance across the difference vectors and zero out components with high variance, focusing on the most consistent rejection-related features.
Step 3: Steering activation. With the identified rejection direction, we intervene in the model’s processing of new queries. We compute the similarity between the query’s representation and the rejection direction. If the similarity exceeds a threshold, indicating a likely conflicting query, we adjust the query’s representation by adding a scaled version of the rejection direction. This steers the query’s representation towards the refusal region of the representation space, encouraging the model to reject inappropriate queries.
6.2 Experiment
@@
To validate the effectiveness of our proposed representation editing method, we conducted comprehensive experiments comparing it with two baseline approaches: Fine-Tuning (FT) and LoRA. We evaluated these methods across various query types and used MT-Bench to assess their impact on general role-playing and conversational abilities. More analysis is presented in the Appendix D.
6.2.1 Baselines
Training details for FT and LoRA are provided in the Appendix C.

### v2

@@ Abstract
Role-Playing Agents (RPAs) have shown remarkable performance in various applications, yet they often struggle to recognize and appropriately respond to hard queries that conflict with their role-play knowledge. To investigate RPAs’ performance when faced with different types of conflicting requests, we develop an evaluation benchmark that includes contextual knowledge conflicting requests, parametric knowledge conflicting requests, and non-conflicting requests to assess RPAs’ ability to identify conflicts and refuse to answer appropriately without over-refusing. Through extensive evaluation, we find that most RPAs behave significant performance gaps toward different conflict requests. To elucidate the reasons, we conduct an in-depth representation-level analysis of RPAs under various conflict scenarios. Our findings reveal the existence of rejection regions and direct response regions within the model’s forwarding representation, and thus influence the RPA’s final response behavior. Therefore, we introduce a lightweight representation editing approach that conveniently shifts conflicting requests to the rejection region, thereby enhancing the model’s refusal accuracy. The extensive experiments validate the effectiveness of our editing method, improving RPAs’ refusal ability of conflicting requests while maintaining their general role-playing capabilities.
Tell Me What You Don’t Know: Enhancing Refusal Capabilities of Role-Playing Agents via Representation Space Analysis and Editing
Wenhao Liu1, Siyu An2, Junru Lu2, Muling Wu1, Tianlong Li1, Xiaohua Wang1, Changze Lv1, Xiaoqing Zheng1thanks: Corresponding author, Di Yin2, Xing Sun2, Xuanjing Huang1 1School of Computer Science, Fudan University  2YouTu Lab, Tencent whliu22@m.fudan.edu.cn, {zhengxq, xjhuang}@fudan.edu.cn {siyuan, junrulu, endymecyyin, winfredsun}@tencent.com
@@ 1 Introduction
Role-Playing Agents(RPAs), ranging from non-player characters in video games Wang et al. (2023) to virtual assistants (Tseng et al., 2024) and interactive educational tools (Wei et al., 2024), are revolutionizing human-computer interaction (Chen et al., 2024b). The growing importance of RPAs in AI applications underscores the need to improve their performance. Previous work in the field of role-playing has primarily focused on enhancing the performance of RPAs through techniques such as prompt-based methods and fine-tuning (Wang et al., 2024a; Zhou et al., 2024; Tu et al., 2023; Li et al., 2023; Chen et al., 2024b; Xu et al., 2024c). To assess these improvements, researchers have introduced several fine-grained evaluation dimensions (Wang et al., 2024b; Chen et al., 2024b, d; Tu et al., 2024; Yuan et al., 2024; Tang et al., 2024; Sadeq et al., 2024), such as assess personality (Wang et al., 2024b) or hallucination (Ahn et al., 2024) of RPAs.
Figure 1: The rejection regions and direct response regions of RPAs in the representation space. The relative distance between a query’s position in the representation space and these regions largely determines whether the model will refuse to answer or respond directly. Different colors in the visualization represent different types of queries, clearly demonstrating the distribution patterns of queries in the representation space and their relationship with knowledge boundaries.
Although these efforts have effectively enhanced the performance of RPAs in terms of role consistency and dialogue capabilities Wang et al. (2024a); Chen et al. (2023), RPAs often struggle when faced with queries that conflict with their role knowledge or capabilities. As a result, they tend to respond directly to queries instead of refusing to answer when faced with such conflicts (Ahn et al., 2024; Sadeq et al., 2024; Tang et al., 2024). For instance, when interacting with an RPA playing the role of Gandalf, if a user queries, “Who murdered Harry Potter’s parents?", an ideal response would be, “I don’t know what you’re talking about. The story of Harry Potter is not part of my world or knowledge." Instead, the RPA might incorrectly reply, “Harry Potter’s parents, James and Lily Potter, were murdered by…" Enhancing the refusal capability of RPAs is crucial for building reliable AI systems. From a safety perspective, it risks the dissemination of inaccurate information. From a user experience standpoint, it compromises role consistency and immersion. From a technical perspective, it indicates limitations in models’ awareness of their knowledge boundaries.
Although some studies have begun to address this issue (Ahn et al., 2024; Sadeq et al., 2024), their scope remains limited, often focusing on specific scenarios such as temporal inconsistencies. There is a lack of systematic research on diverse conflicting scenarios and little exploration of the reasons for RPAs’ performance gap across different types of conflicting queries.
@@
Figure 2: Design of refusal scenarios. Since the knowledge basis for RPAs’ responses typically originates from contextual knowledge and parametric knowledge, we have subdivided the knowledge conflict scenarios into four categories. Among these, the role setting conflict query and role profile conflict query involve conflicts with contextual knowledge, while the factual knowledge conflict query and absent knowledge conflict query involve conflicts with the model’s parametric knowledge. Non-conflict query is used to assess the RPAs’ general role-playing ability.
@@
To answer RQ1 and lay the groundwork for RQ2 and RQ3, we first categorized refusal scenarios based on conflicts with role contextual knowledge and role parametric knowledge, as illustrated in Figure 2. The expected responses from RPAs in these scenarios can range from direct refusal to acknowledging their inability to answer or providing disclaimers about potential errors. To evaluate RPAs’ refusal capabilities, we constructed an evaluation benchmark with queries designed to test various conflict scenarios. We also included non-conflicting queries to assess whether RPAs would excessively refuse to answer. Our evaluation of state-of-the-art models, including GPT-4 and Llama-3, revealed significant differences in their abilities to identify conflicts and refuse to answer across different scenarios. Notably, even advanced models showed unsatisfactory performance when dealing with queries conflicting with role parametric knowledge.
@@
To understand these performance gap, we analyzed model representations under different conflict scenarios (Zou et al., 2023; Liu et al., 2024; Li et al., 2024; Wu et al., 2024). This analysis revealed the existence of rejection regions and direct response regions within the model’s representation space, Figure 1 shows the representation space of Llama3-8B-Instruct when playing Gandalf. Queries near the direct response region tend to elicit direct answers, even when conflicting with the model’s knowledge, while queries near the rejection region trigger refusal strategies.
@@
Based on these findings, we developed a representation editing method to shift conflicting queries from the direct response region toward the rejection region. This approach effectively enhanced the model’s rejection capability while maintaining its general role-playing abilities. Through evaluations using multiple different LLMs as evaluators and human assessment, We compared our method with prompt-based and fine-tuning approaches (Wang et al., 2024a; Zhou et al., 2024; Chen et al., 2023; Li et al., 2023), demonstrating its effectiveness in rejecting conflicting queries without compromising overall performance.
@@ 2 RoleRef: A Benchmark for Evaluating RPA’s Refusal Ability
We first define the refusal capability for RPAs, then introduce the scenarios where RPAs should refuse to answer. Finally, based on the scenarios requiring refusal, we construct our dataset RoleRef (Role-playing agents Refuse to answer). Finally, we propose an evaluation framework to comprehensively measure the role-playing capabilities of RPAs, with a particular emphasis on how they refuse inappropriate or irrelevant questions.
2.1 Refusal Capability
The refusal capability, a critical functionality of RPAs, can be defined as the ability to accurately identify and appropriately reject queries that exceed their knowledge boundaries or conflict with their role settings while maintaining role consistency. This capability encompasses three key dimensions: (1) conflict recognition ability - the capacity to identify conflicts between queries and role knowledge or settings; (2) refusal response ability - providing clear refusal responses with appropriate explanations; and (3) refusal accuracy - avoiding both over-refusal and missed refusals.
@@ 2.2 Scenario Design
  • Role Profile Conflict: The user’s query is in accordance with the role profile, however, it violates specific content within the role profile. For instance, when interacting with an RPA whose role profile states “While Gandalf is powerful, he is not omnipotent." the user asks: “Gandalf, how can I become as omnipotent as you?"
@@
  • Role’s Absent Knowledge Conflict: The character was not present when a specific event occurred. For example, when interacting with an RPA playing the role of Gandalf, the user asks: “Were you there at the moment when Goldberry bid farewell to Frodo and his friends, blessing their journey as they departed on their ponies?".
@@ 2.3 Data Construction
For generating non-conflict queries, we directly prompt the model to generate queries and responses based on the summary content. Additionally, for each query, we require the model to provide the corresponding reference information. The prompts we used are shown in Appendix E.
  Query Type TimeChara RoleRef
Non-conflict $6028$ $11838$
Role Setting - $16455$
Role Profile - $2177$
Factual Knowledge $818$ $12189$
Absent Knowledge $2056$ $2104$
@@ 3 How do existing models perform when facing different types of conflicting queries?
3.1 Models and Metrics
We evaluated these models using the RoleRef dataset. Performance was assessed across 9 dimensions (detailed in Appendix B). Unless otherwise specified, we use GPT-4o as the default evaluator.. Each dimension was scored on a scale of 0 to 2, with the average score reported unless otherwise specified.
@@ 4.1 Analysis via Linear Probes
Previous work has shown that the internal states of LLMs can reveal the model’s knowledge about query truthfulness (Azaria and Mitchell, 2023; Ji et al., 2024). Building on this, we used linear probes to investigate whether models can distinguish between queries that should be refused and those that should be answered. The detailed procedure of probe training is provided in Appendix C.2. The results, shown in Figure 3, reveal following insight:
@@ 4.2 Analysis via t-SNE
Overlap in parametric knowledge conflicts - Direct response region. Representations of most parametric knowledge conflict queries significantly overlap with non-conflict queries. This overlap suggests that these queries within the representation space are positioned in a direct response region, where the model tends to answer directly without recognizing the conflict. For example, when presented with the query “Gandalf, was it you who recommended The Prancing Pony as a safe place to stay for Frodo and his friends?". The representation of this query likely falls within the direct response region, leading to an inappropriate answer. Conversely, for queries whose representations fall further from the non-conflict cluster, the model correctly identifies the false and refuses to answer.
  Methods Non-Conflict Contextual Knowledge Conflict Parametric Knowledge Conflict Average
Role Setting Role Profile Factual Knowledge Absent Knowledge
Prompting $1.94$ $1.94$ $1.62$ $1.16$ $1.02$ $1.53$
FT $1.89$ $1.97$ $1.70$ $1.16$ $1.14$ $1.57$
LoRA $1.84$ $1.97$ $1.61$ $1.22$ $1.14$ $1.56$
Representation Editing $1.92$ $1.96$ $1.78$ $1.19$ $1.02$ $1.57$
Table 5: Human Evaluation Result. We report the average scores across different annotators.
5.1 Representation Editing Method
Step 1: Collecting Activation
For each role, we construct a set of conflict queries and non-conflict queries, represented as:
  • Conflict query set: $\{q_{\text{conflict}}^{i}\}_{i=1}^{N}$
  • Non-conflict query set: $\{q_{\text{non-conflict}}^{i}\}_{i=1}^{N}$
For each query $q$, we obtain the model’s hidden state representation at each layer, denoted as:
  • Conflict query representation at layer $l$: $\mathbf{h}_{\text{conflict}}^{i,l}$
  • Non-conflict query representation at layer $l$: $\mathbf{h}_{\text{non-conflict}}^{i,l}$
where $l=1,2,\dots,L$, and $L$ is the number of layers in the model.
Step 2: Identifying the Rejection Direction
In this step, we calculate the representation differences between conflict and non-conflict queries at each layer to capture the features associated with the model’s refusal behavior.
For each layer $l$, compute the representation difference vector for the $i$-th query pair:
$\Delta\mathbf{h}^{i,l}=\mathbf{h}_{\text{conflict}}^{i,l}-\mathbf{h}_{\text{non-conflict}}^{i,l}$ (1)
Then, calculate the average of all difference vectors to obtain the rejection direction $\mathbf{d}^{l}$ at layer $l$:
$\mathbf{d}^{l}=\frac{1}{N}\sum_{i=1}^{N}\Delta\mathbf{h}^{i,l}$ (2)
To filter out noise and retain features highly related to refusal behavior, we compute the variance for each dimension of the difference vectors. Let $\sigma_{l,j}^{2}$ be the variance of the $j$-th dimension at layer $l$. We zero out dimensions with variance above a threshold $\tau$, resulting in the adjusted rejection direction $\mathbf{d}^{\prime l}$:
$\mathbf{d}_{j}^{\prime l}=\begin{cases}\mathbf{d}_{j}^{l},&\text{if }\sigma_{l,j}^{2}\leq\tau\\ 0,&\text{if }\sigma_{l,j}^{2}>\tau\end{cases}$ (3)
  Methods Non-Conflict Contextual Knowledge Conflict Parametric Knowledge Conflict Average
Role Setting Role Profile Factual Knowledge Absent Knowledge
qwen-max-2025-01-25
Prompt $1.94$ $1.96$ $1.70$ $1.20$ $0.95$ $1.55$
FT $1.91$ $1.99$ $1.73$ $1.23$ $1.15$ $1.60$
LoRA $1.84$ $1.99$ $1.67$ $1.32$ $1.28$ $1.62$
Representation Editing $1.94$ $1.99$ $1.75$ $1.31$ $1.05$ $1.61$
gemini-2.0-pro-exp-02-05
Prompt $1.90$ $1.95$ $1.48$ $0.94$ $0.77$ $1.41$
FT $1.87$ $1.99$ $1.53$ $1.03$ $1.01$ $1.49$
LoRA $1.82$ $2.00$ $1.41$ $1.09$ $1.05$ $1.47$
Representation Editing $1.90$ $1.97$ $1.58$ $1.05$ $0.88$ $1.47$
Doubao-1.5-pro-32k-250115
Prompt $1.97$ $1.92$ $1.50$ $1.23$ $0.94$ $1.51$
FT $1.95$ $1.93$ $1.50$ $1.28$ $1.12$ $1.56$
LoRA $1.89$ $1.93$ $1.45$ $1.31$ $1.16$ $1.55$
Representation Editing $1.97$ $1.95$ $1.56$ $1.27$ $1.00$ $1.55$
Average
Prompt $1.94$ $1.95$ $1.56$ $1.12$ $0.89$ $1.49$
FT $1.91$ $1.97$ $1.58$ $1.18$ $1.10$ $1.55$
LoRA $1.85$ $1.97$ $1.51$ $1.24$ $1.16$ $1.55$
Representation Editing $1.94$ $1.97$ $1.63$ $1.21$ $0.98$ $1.54$
Table 6: Evaluating Llama-3-8B-Instruct Under Different Methods Using Multiple Evaluators.
Step 3: Steering Activation
With the rejection direction for each layer, we intervene in the model’s internal representations when processing new queries.
For a new query $q$, obtain its hidden state representation at layer $l$, $\mathbf{h}^{l}$.
Calculate the similarity between $\mathbf{h}^{l}$ and the rejection direction $\mathbf{d}^{\prime l}$, for example, using cosine similarity:
$\text{sim}(\mathbf{h}^{l},\mathbf{d}^{\prime l})=\frac{\mathbf{h}^{l}\cdot\mathbf{d}^{\prime l}}{\|\mathbf{h}^{l}\|\|\mathbf{d}^{\prime l}\|}$ (4)
If the similarity exceeds a set threshold $\theta$, the query at layer $l$ may require intervention. We add the rejection direction to the original representation proportionally by $\lambda$:
$\mathbf{h}^{l}\leftarrow\mathbf{h}^{l}+\lambda\mathbf{d}^{\prime l}$ (5)
By adjusting the representations at each layer, we gradually guide the model to be more inclined to refuse to answer conflict queries.
5.2 Experiment
To validate the effectiveness of our proposed representation editing method, we conducted comprehensive experiments comparing it with two baseline approaches: Fine-Tuning (FT) and LoRA, details for FT and LoRA are provided in the Appendix C. We evaluated these methods across various query types and used MT-Bench to assess their impact on general role-playing and conversational abilities. More analysis is presented in the Appendix D.